Untitled III

Video & CD-Archive

The ultimate MC, CD, Phono and Video Archive

Version 3.2

Video & CD-Archive

Harald Zottmann

12. September 1995

Contents

1 Author‘s Notes 4 1.1 The Program . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

2 Installation of Untitled III 5 2.1 Installation on a harddisk: . . . . . . . . . . . . . . . . . . . . . . 5 2.2 Installation on a floppy disk: . . . . . . . . . . . . . . . . . . . . 5

3 Program Description 5 3.1 MC - Music Cassettes . . . . . . . . . . . . . . . . . . . . . . . . 5 3.2 CD - Compactdiscs . . . . . . . . . . . . . . . . . . . . . . . . . . 6 3.3 Phono - Longplayers . . . . . . . . . . . . . . . . . . . . . . . . . 6 3.4 Video - Videotapes . . . . . . . . . . . . . . . . . . . . . . . . . . 6

4 File 7 4.1 Load . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 4.2 Save . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 4.3 Save As . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 4.4 Save ASCII . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 4.5 Print . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 4.6 New . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 4.7 Quit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 4.8 Info . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

5 Edit 8 5.1 Cursorcontrol for MC . . . . . . . . . . . . . . . . . . . . . . . . 9 5.2 Cursorcontrol for CD . . . . . . . . . . . . . . . . . . . . . . . . . 9 5.3 Cursorcontrol for Phono . . . . . . . . . . . . . . . . . . . . . . . 9 5.4 Cursorcontrol for Video . . . . . . . . . . . . . . . . . . . . . . . 9 5.5 Cursorcontrol for all modes . . . . . . . . . . . . . . . . . . . . . 10

7 Preferences 12

6 General Keyboardfunctions 11 8 Block Operations 12

8.1 Mark Block . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 8.2 Cut . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 8.3 Copy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 8.4 Erase . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 8.5 Paste . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 8.6 Insert . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 8.7 Convert . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13

9 Layout 14 9.1 Text: Standard . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 9.2 Text: Upper/Lower . . . . . . . . . . . . . . . . . . . . . . . . . . 14 9.3 Adjust: Left/Right/Middle . . . . . . . . . . . . . . . . . . . . . 14

10 Lists 14 10.1 Make . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 10.2 View . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 10.3 Erase . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 10.4 Print . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

11 Ordner 15

12 Extras 16 12.1 Search . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 12.2 F-Keys . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 12.3 Swap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 12.4 Number . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17

13 Tooltypes and Startup Values 17

14 Copyright and Registration 19

15 Future Enhancements 20

1 Author‘s Notes

Untitled III is the name of a program which is able to manage comfortably a database of Music Cassettes (MC), CDs (CD), Longplayers (Phono) and Videocassettes (Video). To avoid misunderstandings, here are a few words about the title of this program. The original name of the program, which was developed and released in 1990, is ‘Untitled‘. In the year 1992, the program was released on the ‘Goldene Serie‘ from DATA Becker with the title ‘Video & CD-Archive‘. This name however says nothing about the programs original main task, the management of MCs and the print-out of tape covers. The Video & CD-Archive is the same program as Untitled V3.0. The current version of this release of Untitled is the version 3.2. There are some major improvements since version 3.0, regarding localization, online help and implementation of other OS2.x features. This program features a very powerful editor with all functions that are well known from real word processors. You have the possibility to print covers for MCs, CDs, Longplayers and Videotapes. You can generate sorted lists, display them on the screen and print them. Those lists may be sorted in different ways. A comfortable search engine can search for every alphanumerical string and guarantees control and fast access to every dataset. There are integrated block functions and many options to convert data between MC, CD, Phono and Video. Functions to calculate spare time and search for spare time keep track on free resources on your videotapes. Because of the numerous data management functions of Untitled III, you will be able to manage your Audio- and Videocollection better than ever before.

1.1 The Program

Untitled III runs on every standard Amiga with at least OS 2.0. Select your printer via the ‘Preferences‘ of the ‘Workbench‘. To make a print-out of tape covers, your printer should be able to print in condensed mode, this means every printer should do. If your Amiga has only 512KByte RAM, please read chapter 13 first. A harddisk is recommended but not necessary anyhow. Untitled III doesn‘t manage its data in uncountable files like other programs do, but in few large files called ‘Ordner‘ (this is a german term). This has many advantages concerning safety of data, access speed and transparency. It leads however to the constraint that you have to record your data in an ascending order. New datasets will be appended to existing datasets The number of a dataset determines its position in the respective ‘Ordner‘. There is a special function implemented in Untitled III which supports this mechanism. To make use of the numerous other functions, you may read the following chapters of this guide file.

2 Installation of Untitled III

2.1 Installation on a harddisk:

The installation of Untitled works with the installer from Commodore. Extract the archive to the RAM disk and click on the installation-icon. Follow the instructions of the installer. The default condition is to store all datafiles and lists in the homedirectory of Untitled III. If you wish to have different pathnames for datafiles and lists, click on the ‘Pref‘ gadget. Now you can enter different pathnames for the ‘Ordner‘ and the program generated lists. Please keep in mind, that the selected directories have to exist because they will not be generated automatically by the program. Leave this menu by clicking the ‘OK‘ gadget. The pathnames for the current mode (MC, CD, Phono or Video) will now be saved. You can repeat these steps for all modes because Untitled supports different pathnames for MC, CD, Phono and Video. This is the end of the installation.

2.2 Installation on a floppy disk:

If you work with floppy disks only, you have to create a workdisk. Just copy all the archive files to a bootdisk with the standard workbench files.

3 Program Description

The following text provides a description of the menu items starting from the left to the right. The menu is hidden in the tile bar of the screen. To activate it, just keep the right mouse button pressed down and move the mousepointer to the top of the screen. Select the desired item and release the button. Most of the menu items can be activated via shortcuts by using the keyboard. The shortcuts are displayed next to the menu items, starting with an inverted ‘A‘, followed by an alphanumerical value. Just hold down the right Amiga Key and press the key which fits to the desired menu item. In addition to the menu functions, there are some other function which can only be invoked via the keyboard and the left Amiga Key. A complete decription in chapter 6 tells more details about that.

3.1 MC - Music Cassettes

The input form for MCs is divided up in Side-A (left part) and Side-B (right part). The upper 15 rows are reserved for the songtitles. The separated two rows are for the band name and the title of the record. The 3 row field at the bottom of the form can be used for additional titles or miscellaneous information like noise reduction, recording date and so on. Please notice that an automatic function for enumerating the MCs writes the tape number in this field, if invoked.

3.2 CD - Compactdiscs

The input form for CDs offers enough space to store the titles of CDs with even very much song titles. There are two times 18 rows for songtitles. The first field in the form takes additional information on the CD. The second row to the left takes the name of the band and the one to the right takes the title of the CD.

3.3 Phono - Longplayers

The input form for longplayers (Phono) takes the additional information in the first row. The following two rows are reserved for the band name and the title of the longplayer. The remaining 17 rows take the song titles.

3.4 Video - Videotapes

The input form of the Video mode differs somehow in comparison to the previous described input forms for MC, CD and Phono. The first row again is for additional information like type and length of the tape. The column ‘Filmtitle‘ should take the titles of the recorded films. This is the place where also the information about artists and director should be located. You should place the title of the film in a separate row and the additional information to that film in the following rows, without a blank row in between.

The total running time of that film and the value of the tapecounter should be located in the same row like the title of the film. Building the list from all tapes, only these films will be added to the list, which have an entry in the ‘Time‘ column. If you don‘t know the running time of the film, please fill scores or something comparable in this column, just to mark this row for the list building function as a valid ‘Filmtitle‘ row. If you own a VCR with a real time counter, you should enter the Start- and the Endtime of the film in the format hh:mm:ss, for example 00:00:00 which means obviously zero, or 01:30:00, which means 1 hour and 30 minutes. The Endtime of the previous film also is the Starttime of the next film. Don‘t forget to fill in the Endtime of the last film of the cassette. Using this data, the program is able to calculate the entries of the ‘Time‘ column and to fill it automatically. You have to press ‘Left Amiga + t‘ to invoke this function. Pressing the key combination ‘Left Amiga + g‘, the program will calculate the spare time on the cassette and fill it in the input form. To use this function, the first row of the input form, with the additional information, has to contain an entry which looks somehow like that: VHS-240 or Beta-180. The important thing about this entry is the length of the tape, measured in minutes. If the column with the time information is filled with the correct length information for each recorded film, with the format h:mm (for example 1:30 which means 1 hour and 30 minutes), the spare time on the cassette will be filled in the first blank row below the last filled row in the filmtitle column.

This information is used by a search function which is able to search for a cassette with a minimum given spare time. Just invoke this function by pressing ‘Left Amiga + h‘ and enter the desired spare time, measured in minutes, in the requester which pops up. The program will search for a cassette which matches this value and will display it. The keys which support this function are desribed in chapter 12.1. The functions which will be invoked with the keyboard are described in chapter 6 once more. If the following description says nothing different, every feature does the same in every mode. The differences between the modes MC, CD, Phono and Video are mentioned at the corresponding places. The modes can be altered also via keyboard:

MC-Mode  Left Amiga + a
CD-Mode  Left Amiga + c
Phono-Mode Left Amiga + p
Video-Mode Left Amiga + v

4 File

4.1 Load

Loads a dataset from the selcted Ordner. Just enter the number of the desired dataset. If you click on the ‘<<‘ gadget, the previous dataset will be loaded, the ‘>>‘ gadget loads the following dataset. The gadget ‘><‘ reloads the current dataset. These functions can be invoked via keyboard also (refer to chapter 6).

4.2 Save

Save the current dataset with the current number to the Ordner. If it is a new dataset, it will be automatically appended at the end of the current Ordner (Destination-Ordner). Please notice that the program distinguishes well between Source- and Destination-Ordner. The Source-Ordner is the one from which a dataset is loaded, or from which the current dataset number is derived. The current Ordner is always the Destination-Ordner. You can check this information at the window title bar of the program. If there is an asterisk displayed after the name of the Ordner, this means that the current dataset is not loaded from the current Destination-Ordner.

4.3 Save As

This function provides the possibility to save a dataset with an user defined number. The number must not be bigger than the number of existing datasets in the current selected Ordner. If you enter 0, the dataset will be appended at the current Ordner (Destination-Ordner).

4.4 Save ASCII

The current dataset will be saved to an ASCII file. You now can load this file into any DTP program just like PageStream. This allows you to modify the layout and font, and print it in graphics mode.

4.5 Print

A requester pops up and lets you select the numbers of the datasets to print. If you don‘t select any numbers and click on the ‘Print‘ gadget, the current dataset will be printed. In addition to that, you can select some attributes for the print-out. These attributes are valid for the whole print session. If you want to print more than one dataset, adjust the paper of the connected printer. Untitled III automatically makes a formfeed at the end of each printed page.

4.6 New

The current dataset (Mask-1 or Mask-2) will be erased from memory. The contents of the internal clipboard remains untouched.

4.7 Quit

Quit the program. You can quit Untitled III by clicking on the ‘Esc‘ gadget too. If the requesters aren‘t switched off (default), a requester pops up just to make sure. If you select ‘Yes‘, Untitled III checks if there are any unsaved changes. Untitled III will automatically switch to the mode with unsaved changes and give you the possibility to abort this command.

4.8 Info

Gives you some information about Untitled III.

5 Edit

The following terms are equal for MC, CD, Phono and Video:

Arrow left  4 NUM
Arrow right 6 NUM
Arrow up  8 NUM
Arrow down 2 NUM

and vice versa. The suffix NUM means the keys are to be found on the number pad.

5.1 Cursorcontrol for MC

Position  Key

First Title (A)

7 NUM

First Title (B)

9 NUM

Bandname (A)

1 NUM

Bandname (B)

3 NUM

Tapenumber 5 NUM

Swap A-Side / B-Side

Tab

Swap B-Side / A-Side

Tab

5.2 Cursorcontrol for CD

Position  Key
Bandname  7 NUM
CD-Title  9 NUM

First Title (left)

1 NUM

First Title (right)

3 NUM

Info row 5 NUM

Swap Left / Right

Tab

Swap Right / Left

Tab

5.3 Cursorcontrol for Phono

Position  Key
First Title  1,3 NUM
Bandname  7,9 NUM
Info row  5 NUM

5.4 Cursorcontrol for Video

Position  Key

First Filmtitle

1,7 NUM

First Counter

3,9 NUM

Info row 5 NUM

Swap Left / Right

Tab

Swap Right / Left Shift + Tab

5.5 Cursorcontrol for all modes

Action  Key
Prev Word  Ctrl + Arrow left
Next Word  Ctrl + Arrow right
Beg of line  Shift + Arrow left
End of line  Shift + Arrow right
Beg of input form Alt + Arrow left
End of input form Alt + Arrow right
Three rows up  Ctrl + Arrow up
Three rows down Ctrl + Arrow down
First row  Shift (Alt) + Arrow up
Last row  Shift (Alt) + Arrow down
Insert mode  0 NUM
Overwrite mode  . NUM

Cursorcontrol can be done via mouse too. The ‘Delete‘ key will erase the character to the right of the cursor. If there are no more characters to the right of the cursor, the following rows will move one row up. The ‘Backspace‘ key will erase the character to the left of the cursor. The following text will move to the left too. The ‘Return‘ key moves the cursor and the text to the right of the cursor to the beginning of the next row, if the last row of the form is blank. If not, the cursor will only be repositioned to the next row.

6 General Keyboardfunctions

Action  Key
Load first dataset Left Amiga + 7 NUM
Load last dataset  Left Amiga + 1 NUM
Ten datasets back Left Amiga + 3 NUM
Ten datasets ahead Left Amiga + 9 NUM
‘<<‘ gadget  Left Amiga + 4 NUM
‘<<‘ gadget  Left Amiga + F1
‘><‘ gadget  Left Amiga + 5 NUM
‘><‘ gadget  Left Amiga + F2
‘>>‘ gadget  Left Amiga + 6 NUM
‘>>‘ gadget  Left Amiga + F3
Undo  Left Amiga + u
Undo  Left Amiga + F4
Redo  Left Amiga + r
Redo  Left Amiga + F5
‘A‘ gadget  Left Amiga + F6
‘B‘ gadget  Left Amiga + F7
‘1‘ gadget  Left Amiga + F8
‘2‘ gadget  Left Amiga + F9
Pref  Left Amiga + F10

Undo / Redo cand handle between 2 and 20 steps (refer to chapter 13). If the system is low on memory, Untitled III switches to 2 steps automatically at program startup. Both functios can be invoked with the gadgets of the same name too. With the gadgets ‘A‘ and ‘B‘ you can mark the whole A-Side or B-Side of a tape in the MC mode. With the gadgets ‘1‘ and ‘2‘ you can switch between two different datasets of one mode. Together with the block functions, you can easily copy data from one dataset to another. Funktions which can only be invoked via keyboard:

Action  Key
Copy block and erase it  Left Amiga + b
Delete line  Left Amiga + k
Insert current Systemdate Left Amiga + d
Insert current Systemtime Left Amiga + z
Erase list from memory  Left Amiga + l

Only in Video mode available:

Action  Key
Init ‘Time‘ column  Left Amiga + t
Calculate spare time  Left Amiga + g
Search for given spare time Left Amiga + h

7 Preferences

A window pops up with some information about the current status and global information of the program.

1. Pathnames for Ordners and Lists. These pathnames can be modified right here.

2. Number of entries in the different Ordners, belonging to the current mode. Existence of Lists.

3. Source-Ordner, Destination-Ordner and number of the current datasets in Mask-1 and Mask-2. If the colour of a number of a certain Mask differs from the rest of the text, the dataset in this Mask is modified and not saved yet.

To validate changes at the pathnames, you have to click on the ‘OK‘ gadget. The pathnames will be saved in the file ‘VCA.CFG‘. This file is automatically loaded at every program startup.

8 Block Operations

8.1 Mark Block

Marks a block, starting at the current position of the cursor. To size the block, you can use either the cursor keys or the mouse. The marked text will be displayed in a different colour. The mark block function can be disabled by clicking on the ‘Esc‘ gadget or by pressing Escape on the keyboard. This function can also be enabled by double clicking on a certain character The starting point of the marked block will also be displayed in a different colour. If you click again on this character, you can disable the mark block function alternatively.

8.2 Cut

Cuts the marked block. To use this function, the marked block must not cross the borders of a certain field. The block will be copied in the internal clipboard and can be pasted or inserted somewhere else.

8.3 Copy

The marked block will be copied in the internal clipboard. The contents of the current dataset remains untouched. The block borders are not limited in any way.

8.4 Erase

The marked block will be erased from the dataset. The contents of the internal clipboard remains untouched.

8.5 Paste

The contents of the internal clipboard will be pasted at the current cursor position of the displayed dataset.

8.6 Insert

The contents of the internal clipboard will be inserted at the current cursor position of the displayed dataset. This operation does not change the contents of other fields in the mask, except the one with the cursor.

8.7 Convert

This function provides a fast and easy data exchange mechanism between the different modes MC, CD and Phono. If you invoke this function for example in the CD mode, you can select whether you want to convert the data in the current dataset from the CD format to the MC or Phono format. For example, if you want to record a few of your CDs for a friend, you can easily build and print the MC covers without a single keystroke.

Example: If you want to test this out, you have to load or write a CD dataset first. Then you have to invoke the ‘Convert‘ function from the menu or via shortcut. If the CD has not more than 15 title rows, you can convert the dataset to a single MC side. (gadget ‘CD − > A‘). Otherwise you click on the ‘CD − > AB‘ gadget to convert the whole CD dataset. Don‘t panic, you‘re CD dataset remains untouched by this operation. Now swap to the MC mode via menu or shortcut and select a blank input form. Position the cursor to the left upper edge of the input form and choose the function ‘Paste‘ from the ‘Block‘ menu. The converted dataset will now be copied to the MC input form. You now can save, print or work with this MC dataset. If you want to convert another CD for the B side of the MC, just repeat the previous steps. The only difference is, that you have to position the cursor at the upper left edge of the B-Side of the MC input form, before you select the ‘Paste‘ command from the menu.

Hint: If you convert datasets where the source dataset has more titles than the destination dataset, you have to take care that no data will get lost. For example, a CD may have (2 x 18) 36 rows for song titles, whereas one side of a MC has only 15 rows for song titles.

9 Layout

9.1 Text: Standard

This is a block operation. If you invoke this function without a block being currently marked, the function works on the whole dataset. The text will be changed to lower case whereas every first charcter of a new word will be changed to upper case.

9.2 Text: Upper/Lower

Works like the previous function, but changes the text to upper or lower case.

9.3 Adjust: Left/Right/Middle

This is a block function also. If no block is marked, it works on the line where the cursor is currently positioned. The text in this line will be adjusted to the left border, the right border or be centered.

10 Lists

10.1 Make

This function can build lists of the current Destination-Ordner and sort them. The contents of the list is different for the modes MC, CD, Phono and Video:

MC:  Bandname Title Dataset number A
Bandname Title Dataset number B
CD:  Bandname Title Dataset number Info row
Phono: Bandname Title Dataset number Info row
Video: Filmtitle  Time Dataset number Info row

If the gadget ‘New‘ is pressed, a fresh list will be build from the data of the current Destination-Ordner. The list is sorted with ascending dataset numbers. If the gadget ‘Article‘ is pressed in the modes MC, CD or Phono, the articles (the, der, die or das) will be appended to the bandname if they occur in front of the bandname. This function supports the sorting abilities and should be enabled everytime you build a new list. The gadget ‘Empty‘ means that blank datasets will be ignored by the list building function. If the gadget ‘Sort‘ is pressed additionaly, the list will be sorted in an alphabetically ascending way. If an unsorted, up to date list already exists, it can optionally be sorted without rebuilding it, if the gadget ‘New‘ is disabled. The gadget ‘UC=lc‘ means that the program will not distinguish between upper and lower case when sorting the list. By invoking the ‘Info‘ gadget, you can treat the program to sort the list according to the Info row. Otherwise it will be sorted according to the bandnames (MC, CD and Phono) or the Filmtitles (Video).

10.2 View

This function displays the list, according to the current Destination-Ordner, on the screen. You can move the list up and down with the slider gadget on the right border of the list or with the cursor keys. In the CD, Phono or Video mode, you can move the list horizontally with the slider gadget at the bottom of the list. If the list is sorted alphabetically, you can press a key (a-z) or (0-9) to jump to the corresponding point in the list. If you click once on an entry of the list, it will be displayed inverted. Clicking on the same entry for a second time will load this dataset and you will get back to the edit mode. Please notice, that the list which was displayed at last will remain in memory, until you want to view a different list. This feature increases the speed of this function dramatically, especially if you are working with floppys. If your system runs low on memory, you can flush the list from memory by pressing ‘Left Amiga + l‘.

10.3 Erase

This function deletes the list corresponding to the current Destination-Ordner, from harddisk or floppy.

10.4 Print

This function lets you print a list with page numbering an automatic formfeed. If the ‘Date‘ gadget is enabled, the current system date will be printed on every page. With the ‘Info‘ gadget, you can select whether the Info row is being printed, in the CD, Phono or Video mode, or not.

11 Ordner

You have the choice to select one of five Ordners in every mode, where you can store and load your datasets from. If you don‘t want to work with more than one Ordner, you should only work with the default Ordner ‘Ordner 1‘. There is no limitation concerning the number of datasets, if there is enough space on floppy or harddisk. On a formatted 880 KByte disk, you can store about 650 datasets of the MC or CD type, 1300 datasets of the type Phono or 1050 datasets of the type Video. If there are more datasets than this, a harddisk is recommended but not necessary, if you split the data to different Ordners. The filenames for Ordner and lists are:

MC:  MC Ordner 1 - MC Ordner 5

MC Ordner 1.Liste - MC Ordner 5.Liste

CD:  CD Ordner 1 - CD Ordner 5

CD Ordner 1.Liste - CD Ordner 5.Liste

Phono: Phono Ordner 1 - Phono Ordner 5

Phono Ordner 1.Liste - Phono Ordner 5.Liste

Video: Video Ordner 1 - Video Ordner 5

Video Ordner 1.Liste - Video Ordner 5.Liste

12 Extras

12.1 Search

This is a very comfortable search function which searchs in the current Destination-Ordner of the current mode (refer to window title bar). Enter a word or a whole sentence to search for in the string gadget named ‘Term to search for:‘. You can select the dataset number where to begin with the search, by using the slider gadget. If you select the highest dataset number with the slider gadget, the search will begin at the end of the Ordner and go down to the start direction. Otherwise the search goes from the start up to the end. If the gadget ‘UC=lc‘ is activated (default), there will be no difference between upper and lower cases. This means that using the search term ‘hello‘, the function will find the terms ‘Hello‘ and ‘HELLO‘ equally.

The gadget ‘Words‘ means that the search term must not be a part of another bigger word. Otherwise the search term ‘computer‘ will be found in the word ‘Personalcomputer‘ too. The ‘Title‘ gadget means that the program will only search in the bandname and Recordtitle field in the MC, CD amd Phono mode and in the Info row of the Video mode. Otherwise, the funtion scans the whole dataset. Start the search function with the ‘Search‘ gadget. There will be a fuel gauge which displays the part of the Ordner, where the search function currently works. The window title bar displays the direction of the search mode. You can alter the direction at any time with the ‘<<‘ and ‘>>‘ gadgets. Alternatively you can use the cursor keys. If the term is found, the search will be interrupted and the corresponding dataset will be displayed, with the found search term being marked. You can continue the search with the previous mentioned gadgets, or you can load the displayed dataset with the ‘><‘ gadget. Otherwise you can cancel the search with Escape. If you cancel the search, the original dataset will be displayed.

12.2 F-Keys

This function pops up a window which displays the meaning of the F-Keys. Invoking the ‘Save‘ gadget, you can save the F-Keys to the file ‘VCA.FNK‘, or you can load the F-Keys from disk using ‘Load‘. In addition to that, you can send the meaning of the F-Keys to a printer. To record something to a F-Key buffer, you have to do the following. Mark a one line block with the function ‘Mark‘. This block must not cross the borders of a field. Now press the F-Key (F1 - F10) which shall take the marked block. That‘s all for now. If you press that F-Key in the edit mode, the contents of the F-Key will be inserted in the input form at the cursor position. Note that this function distinguishes between the Insert and the Overwrite mode.

12.3 Swap

Swaps the A-Side and the B-Side of the current dataset. This function is only available in the MC mode.

12.4 Number

Inserts the current number of the dataset in the input form. If the number was not defined so far, the next available number of the Ordner will be selected.

13 Tooltypes and Startup Values

To configure the program upon startup, you can select a few startup values, in the infofile of Untitled III. Click once on the program icon and select the ‘Information‘ menu from the workbench. You can now select the following ‘Tooltypes‘.

Possible values:

        MODUS=x         x is a number value between 1 and 4.
                        This means:     1       MC-Mode
                                        2       CD-Mode
                                        3       Phono-Mode
                                        4       Video-Mode
                        The selected mode will be activated at
                        startup.
        ORDNER=x        x is a number value between 1 and 5.
                        This means:     1       Ordner_1
                                        2       Ordner_2
                                        3       Ordner_3
                                        4       Ordner_4
                                        5       Ordner_5

UNDO=x          x is a number value between 2 and 20.
It means the maximum number of Undo /
Redo steps. Please notice that this
value is not too high if you don`t
have enough memory. If the program
can`t get enough memory for the desired
value, it will allocate memory for 2
steps only.
REQUESTER=x     This means:     x = 0   Requester off
                x = 1   Requester on
INSERT=x        This means:     x = 0   Overwrite
                x = 1   Insert
ZOOM=x          This means      x = 0   Zoom off
                x = 1   Zoom on
LANGUAGE=x      This means      x = 0   German
                x = 1   English
STYLE=x         This means      x = 0   Old Requester
                x = 1   New Requester

If your Amiga has only 512Kbyte of memory, you should select UNDO with the option x=2, in order to save memory.

14 Copyright and Registration

This program is Shareware. If you like it and use it, please register using the orderform that comes with this package.

                        Harald Zottmann
                        Friedenstrasse 27
                        63743 Aschaffenburg
                        or
                        Harald Zottmann
                        Am Altenbruch 43
                        40822 Mettmann
            you can reach me on the Internet via E-Mail:
                  harald.zottmann@ntc.nokia.com

If you register, you will get a Keyfile which switches off that boring registration delay requester. This Keyfile enables you to use all future releases of the program. Furthermore you will get the newest version of the program, optimized for your micro processor (please note your processor type on the orderform).

This program comes without any warranty. I don‘t accept responsibility to anyone for the consequences of using it or for whether it serves any particular purpose or works at all. Bugreports and suggestions are welcome.

15 Future Enhancements

Planned Enhancements:

1. Importmodules and Exportmodules for other well known programs 2. Printeroutput additionaly in LaTEX or Postscript

3. Improved GUI

4. Improved Clipboard support

5. User selectable fonts

6. Powerpacker support

7. Sizing of list view window

8. Quicksort for lists

9. and much more...

Updates will only be published via the Aminet or other PD sources.